# Water wave scattering by a surface-mounted rectangular anisotropic elastic plate

### v2

@@ 5 Radiation potentials
where, in abuse of notation, we have written $w_{j}(\mathbf{x^{\prime}})=w_{j}([x^{\prime},y^{\prime},0]^{\intercal})=w_{j}(x^{\prime},y^{\prime})$. Applying the constant panel discretisation method used in (3.3) then yields the matrix equation
@@ 6 Scattering by an elastic plate
The solution to the coupled fluid-plate system of equations (14) is obtained by expanding in the dry modes of vibration of the plate, following the theory of newman1994wave. That is, the total potential is taken to be a superposition of the incident potential, the diffraction potential plus a weighted sum of the radiation potentials
@@
Note that in this paper $N=30$ plate modes were considered, which was considered sufficient based on the energy balance validation in the following section. Also note that our implementation of this method has been validated using code developed by meylan2002wave for the case of a square isotropic plate with free edges.
@@ 7 Energy balance calculations
such that the corresponding free surface elevation is $\eta^{\rm in}(x,y)=\mathrm{e}^{\mathrm{i}kx}$. Let $(r,\theta,z)$ be the system of cylindrical coordinates with $x=r\cos\theta$ and $y=r\sin\theta$. For large $r$, the potential can be approximated as (Mei2005)
@@
Suppose the source point of the Green’s function is shifted to $(x^{\prime},y^{\prime})$. Let $d(x^{\prime},y^{\prime})$ and $\alpha(x^{\prime},y^{\prime})$, respectively, be the radial and angular coordinates of the point $(x^{\prime},y^{\prime})$ with respect to the global origin $(0,0)$. Then as $r\to\infty$, the Green’s function is asymptotically given by (Falnes2002)
@@
Conservation of energy implies that the optical theorem holds, which is (Mei2005)
@@ 8 Results
The next case we consider is a square anisotropic plate, as parametrised by 10.1115/1.4053090. These parameters, which are reproduced in Table 1, describe a graphite/epoxy laminated plate in which the tensor describing the material is rotated by 45 in the $xy$-plane. As such, the modes of vibration in Figure 2 are symmetric or antisymmetric with respect to the lines $y=x$ and $x+y=a$. The response of the surface-mounted square anisotropic plate to a plane wave with incident angle $\chi=0$ is considered in Figures 911. We observe three resonant peaks in the kinetic energy spectrum (Figure 9), which are predominantly associated with excitation of the first three modes of vibration in order of increasing frequency. The second and third modes of vibration are not degenerate, and neither are antisymmetric with respect to the line $y=b/2$, so both can be excited. We remark that the far field patterns (Figure 10b, d and f) and the surface elevations (Figure 11) do not possess any symmetry. Observe that while Figures 10b and 11a may appear at first glance to suggest a symmetric response, this is due to the first mode shape (Figure 2i) being close to symmetric—there are in fact subtle asymmetries in Figure 11.
@@
Lastly, we consider anisotropic plates with free edges, in both square ($a=b=1$ m) and rectangular ($a=2b=2$ m) geometries. Modes of vibration are illustrated in Figure 22, which also includes the case of a square isotropic plate for the purpose of validation against LEISSA1973257. Kinetic energy spectrums for square and rectangular anisotropic plates with free edges are given in Figures 23 and 26, respectively. Each of these has only one resonant peak in the frequency interval considered, which is noticeably broader than peaks in the corresponding spectra for clamped plates (Figures 9 and 19, respectively). Figures 24 and 27 show the dry mode coefficient amplitudes and far field patterns at the resonant frequency. We observe that significant excitation occurs over many modes. In contrast, corresponding figures for clamped plates (Figures 10 and 20) show that the resonances are predominantly due to excitation of a single mode. Figures 25 and 28 show the surface elevations for square and rectangular anisotropic plates with free edges, respectively.
@@ 9 Conclusion
The main motivation of this work was an application to the modelling of PWECs, and this remains as an area of future work. The method presented here can be used (essentially verbatim) for a surface mounted rectangular piezoelectric plate, with the rigidity coefficients $D_{ij}$ becoming complex. The only necessary adjustment resulting from this would be to the dry mode expansion method presented in §6, because the set of plate eigenmodes $w_{j}$ may no longer be orthogonal in general, which would cause the hydrodynamic matrices $K_{\rm stiff}$, $M$ and $C$ to have non-zero off diagonal entries proportional to $\iint_{\Gamma\times\{0\}}w_{j}w_{m}^{*}\,\mathrm{d}S$. With this detail taken into account, we stress that our approach will allow the anisotropy of the piezoelectric material within the WEC to be modelled appropriately in three dimensions. We note that PWECs have not been considered in this paper because the determination of the complex rigidity coefficients from the poling direction, circuiting conditions, and layering of the piezoelectric material is a complicated modelling problem (renzi2016hydroelectromechanical). Another possible extension of this work is to arbitrary shaped plates, as considered by meylan2002wave for isotropic plates with free edges. Although the modes of vibration would need to be calculated using a different method to the one presented in §4 (e.g., the finite element method could be used), the hydrodynamic component of the solution could remain unchanged. Lastly, we mention the problem of extending this study to submerged horizontal plates. This case is more complicated from a mathematical perspective because it leads to hypersingular integral equations.
@@ Appendix A Derivation of the Green’s function
so we compute
$A(\mu)$ $=-\frac{1}{2\pi}\frac{1}{\mu\sinh\mu H-\alpha\cosh\mu H}$ (63)
@@
Note that we must integrate beneath the pole of the second integrand in order to satisfy the Sommerfeld radiation condition (57d). After expressing $\cosh\mu(z+H)=\tfrac{1}{2}(\mathrm{e}^{\mu(z+H)}+\mathrm{e}^{-\mu(z+H)})$ and using the fact that (Bracewell1999)
@@ Appendix B Beam eigenfunctions
In the case that the beam is free at both ends, as in equation (38), the ansatz (67) gives rise to the same transcendental equation for the beam eigenvalues $\kappa_{m}$ as the clamped case, namely

### v3

@@ 5 Radiation potentials
where, in abuse of notation, we have written $\mathrm{w}_{j}(\mathbf{x^{\prime}})=\mathrm{w}_{j}([x^{\prime},y^{\prime},0]^{\intercal})=\mathrm{w}_{j}(x^{\prime},y^{\prime})$. Applying the constant panel discretisation method used in (3.3) then yields the matrix equation
@@ 6 Scattering by an elastic plate
The solution to the coupled fluid-plate system of equations (21) is obtained by expanding in the dry modes of vibration of the plate, following the theory of Newman (1994). That is, the total potential is taken to be a superposition of the incident potential, the diffraction potential plus a weighted sum of the radiation potentials
@@
Note that in this paper $N=30$ plate modes were considered, which was considered sufficient based on the energy balance validation in the following section. Also note that our implementation of this method has been validated using code developed by Meylan (2002) for the case of a square isotropic plate with free edges.
@@ 7 Energy balance calculations
such that the corresponding free surface elevation is $\eta^{\rm in}(x,y)=\mathrm{e}^{\mathrm{i}kx}$. Let $(r,\theta,z)$ be the system of cylindrical coordinates with $x=r\cos\theta$ and $y=r\sin\theta$. For large $r$, the potential can be approximated as (Mei et al., 2005)
@@
Suppose the source point of the Green’s function is shifted to $(x^{\prime},y^{\prime})$. Let $d(x^{\prime},y^{\prime})$ and $\alpha(x^{\prime},y^{\prime})$, respectively, be the radial and angular coordinates of the point $(x^{\prime},y^{\prime})$ with respect to the global origin $(0,0)$. Then as $r\to\infty$, the Green’s function is asymptotically given by (Falnes, 2002)
@@
Conservation of energy implies that the optical theorem holds, which is (Mei et al., 2005)
@@ 8 Results
The next case we consider is a square anisotropic plate, as parametrised by An et al. (2021). These parameters, which are reproduced in Table 1, describe a graphite/epoxy laminated plate in which the tensor describing the material is rotated by 45 in the $xy$-plane. As such, the modes of vibration in Figure 2 are symmetric or antisymmetric with respect to the lines $y=x$ and $x+y=a$. The response of the surface-mounted square anisotropic plate to a plane wave with incident angle $\chi=0$ is considered in Figures 911. We observe three resonant peaks in the kinetic energy spectrum (Figure 9), which are predominantly associated with excitation of the first three modes of vibration in order of increasing frequency. The second and third modes of vibration are not degenerate, and neither are antisymmetric with respect to the line $y=b/2$, so both can be excited. We remark that the far field patterns (Figure 10b, d and f) and the surface elevations (Figure 11) do not possess any symmetry. Observe that while Figures 10b and 11a may appear at first glance to suggest a symmetric response, this is due to the first mode shape (Figure 2i) being close to symmetric—there are in fact subtle asymmetries in Figure 11.
@@
Next, we consider anisotropic plates with free edges, in both square ($a=b=1$ m) and rectangular ($a=2b=2$ m) geometries. Modes of vibration are illustrated in Figure 22, which also includes the case of a square isotropic plate for the purpose of validation against Leissa (1973, Table C15). Kinetic energy spectrums for square and rectangular anisotropic plates with free edges are given in Figures 23 and 26, respectively. Each of these has only one resonant peak in the frequency interval considered, which is noticeably broader than peaks in the corresponding spectra for clamped plates (Figures 9 and 19, respectively). Figures 24 and 27 show the dry mode coefficient amplitudes and far field patterns at the resonant frequency. We observe that significant excitation occurs over many modes. In contrast, corresponding figures for clamped plates (Figures 10 and 20) show that the resonances are predominantly due to excitation of a single mode. Figures 25 and 28 show the surface elevations for square and rectangular anisotropic plates with free edges, respectively.
@@
Finally, in Figure 29 we present kinetic energy spectra for plates with simply supported edges. While this is primarily included to serve as a benchmark calculation, we note that the resonant frequencies are lower than those of clamped plates (e.g., compare Figures 3 and 29 for isotropic plates), owing to their lower frequencies of vibration (Leissa, 1973).
Figure 29: Kinetic energy spectra for isotropic, orthotropic and anisotropic square plates ($a=b=1$ m) with simply-supported edges. Only the resonances of the isotropic case are marked.
@@ 9 Conclusion
The main motivation of this work was an application to the modelling of PWECs, and this remains as an area of future work. The method presented here can be used (essentially verbatim) for a surface mounted rectangular piezoelectric plate, with the rigidity coefficients $D_{ij}$ becoming complex. The only necessary adjustment resulting from this would be to the dry mode expansion method presented in §6, because the set of plate eigenmodes $w_{j}$ may no longer be orthogonal in general, which would cause the hydrodynamic matrices $K_{\rm stiff}$, $M$ and $C$ to have non-zero off diagonal entries proportional to $\iint_{\Gamma\times\{0\}}w_{j}w_{m}^{*}\,\mathrm{d}S$. With this detail taken into account, we stress that our approach will allow the anisotropy of the piezoelectric material within the WEC to be modelled appropriately in three dimensions. We note that PWECs have not been considered in this paper because the determination of the complex rigidity coefficients from the poling direction, circuiting conditions, and layering of the piezoelectric material is a complicated modelling problem (Renzi, 2016). Another possible extension of this work is to arbitrary shaped plates, as considered by Meylan (2002) for isotropic plates with free edges. Although the modes of vibration would need to be calculated using a different method to the one presented in §4 (e.g., the finite element method could be used), the hydrodynamic component of the solution could remain unchanged. Lastly, we mention the problem of extending this study to submerged horizontal plates. This case is more complicated from a mathematical perspective because it leads to hypersingular integral equations.
@@ Appendix A Derivation of the Green’s function
for z=0z=0¡ so we compute
$A(\mu)$ $=-\frac{1}{2\pi}\frac{1}{\mu\sinh\mu H-\alpha\cosh\mu H}$
@@
Note that we must integrate beneath the pole of the second integrand in order to satisfy the Sommerfeld radiation condition (65d). After expressing $\cosh\mu(z+H)=\tfrac{1}{2}(\mathrm{e}^{\mu(z+H)}+\mathrm{e}^{-\mu(z+H)})$ and using the fact that (Bracewell, 2000)
@@ Appendix B Beam eigenfunctions
In the case that the beam is free at both ends, as in equation (45), the ansatz (74), upon formation of a matrix analogous to that in (75), gives rise to the same transcendental equation for the beam eigenvalues $\kappa_{m}$ as the clamped case, namely
@@
The eigenvalues and eigenmodes for the simply supported case (46) have the following simple form:
$\kappa_{n}$ $=n\pi,$ (79a)
$u_{n}(x)$ $=\sqrt{2}\sin(n\pi).$ (79b)